2) Shining Star Company uses an automated process to clean and polish its souvenir items. For March, the company had the following activities: Required: Prepare a production cost worksheet using the FIFO method.

Beginning work in process inventory

3,000 items, 1/3 complete with regards to conversion costs

Units placed in production

12,000 units

Units completed

9,000 units

Ending work in process inventory

6,000 items, 2/5 complete with regards to conversion costs

Cost of beginning work in process

$2,500

Direct material costs, current

$9,000

Conversion costs, current

$8,320

Direct materials are placed into production at the beginning of the process and conversion costs are incurred evenly throughout the process.

Assignment of costs
Work in process, beginning
Completion of beginning
Total beginning inventory
Started and Completed
Total costs transferred out $13,400
Work in process, ending
   Direct materials
   Conversion
Costs accounted for $19,820
Answer #1

Direct materials are placed into production at the beginning of the process and conversion costs are incurred evenly throughout the process.

Assignment of costs
Work in process, beginning 2500
Completion of beginning (2000*.80) 1600
Total beginning inventory 4100
Started and Completed 9300
Total costs transferred out $13,400
Work in process, ending
   Direct materials 4500
   Conversion 1920 6420
Costs accounted for $19,820

Direct material equivalent unit = 6000+6000 = 12000 Cost per equivalent unit of direct material = 9000/12000 = 0.75

Conversion equivalent unit = (3000*2/3+6000*100%+6000*2/5) = 10400 cost per equivalent unit = 8320/10400 = 0.80